# Prompts Are Programs Too! Understanding How Developers Build Software Containing Prompts

### v1

@@ Abstract.
The introduction of generative pre-trained models, like GPT-4, has introduced a phenomenon known as prompt engineering, whereby model users repeatedly write and revise prompts while trying to achieve a task. Using these AI models for intelligent features in software applications require using APIs that are controlled through developer-written prompts. These prompts have powered AI experiences in popular software products, potentially reaching millions of users. Despite the growing impact of prompt-powered software, little is known about its development process and its relationship to programming. In this work, we argue that some forms of prompts are programs, and that the development of prompts is a distinct phenomenon in programming. We refer to this phenomenon as prompt programming. To this end, we develop an understanding of prompt programming using Straussian grounded theory through interviews with 20 developers engaged in prompt development across a variety of contexts, models, domains, and prompt complexities.
Through this study, we contribute 14 observations about prompt programming. For example, rather than building mental models of code, prompt programmers develop mental models of the FM’s behavior on the prompt and its unique qualities by interacting with the model. While prior research has shown that experts have well-formed mental models, we find that prompt programmers who have developed dozens of prompts, each with many iterations, still struggle to develop reliable mental models. This contributes to a rapid and unsystematic development process. Taken together, our observations indicate that prompt programming is significantly different from traditional software development, motivating the creation of tools to support prompt programming. Our findings have implications for software engineering practitioners, educators, and researchers.
@@ 1. Introduction
The introduction of generative pre-trained models (e.g., GPT-4 (Achiam et al., 2023), Dall-E (Betker et al., 2023))—also known as foundation models (FMs)—have drastically changed how software is built by developers. AI programming assistants that generate code (e.g., GitHub Copilot (Chen et al., 2021)) have aided developers across a variety of tasks such as writing significant portions of code, learning new APIs and programming languages, writing tests (Liang et al., 2024), and information seeking behaviors (Haque et al., 2024) thereby improving the productivity of programmers (Ziegler et al., 2022; Peng et al., 2023; Ziegler et al., 2024). Recently, instruction-tuned (Ouyang et al., 2022) large language models (LLMs), such as ChatGPT (ope, 2024a), have assisted developers with a broader range of software development-related tasks via writing natural language prompts. This includes resolving code issues, developing new features, refactoring, and writing configuration files (Chavan et al., 2024).
@@
The rising prominence of prompts has ushered in a new phenomenon known as prompt engineering, whereby model users repeatedly write and revise natural language prompts to achieve a task. These prompts enable more intelligent features when integrated in popular code-based software applications (Parnin et al., 2023), such as Google Search (goo, 2024) and Microsoft Office (mic, 2024b), reaching potentially millions of users (mic, 2024a). As of January 2024, engineered prompts have also underpinned the creation of over 3 million custom versions of ChatGPT for a specific task, known as GPTs (ope, 2024b).
@@
Despite the budding impact of prompt-powered software, little is known about the prompt engineering process and its relationship to programming. To the best of our knowledge, two such studies in software engineering currently exist. One is Dolata et al. (2024)’s study of 52 freelance developers on the challenges of developing solutions based on generative AI. Their interview focuses on the positive and negative experiences with generative AI, project uncertainties, and views on freelancing. The other is Parnin et al. (2023)’s study of 12 professional software developers integrating generative AI into products. Their interview topics include the participant’s motivation for using AI in the product, major tasks for building the generative AI application, prompt engineering, testing, tooling, challenges, learning related skills, and concerns with AI. While these studies offer interesting preliminary insights on prompt development, they are constrained to solely freelance and professional software development, limiting the generalizability of the results to other types of programming. Further, these studies do not place targeted focus on the process of developing prompts in relation to programming activities. We address the gaps in this literature by following a more systematic and rigorous qualitative methodology—known as Straussian grounded theory (Corbin and Strauss, 2015)—and performing maximum variation sampling (Suri, 2011) to recruit a diverse sample of programmers with varying programming contexts, models, roles, application domains, and prompt complexities to understand the process of using prompts as part of creating software applications.
In this work, we argue that some forms of prompts are programs. We view the development of such prompts and other natural language-powered software as a distinct phenomenon in programming. We refer to it as prompt programming, following prior work (Jiang et al., 2022). We note that the concept of programming using natural language is not novel; in 1979, Edsger Dijkstra discussed a phenomenon where one could “instruct [a machine] in our native tongues” (Dijkstra, 2005). This vision has only recently been realized, with the advent of FMs providing the means for technical feasibility and widespread use of such programs.
@@
We argue that prompt programming is a phenomenon that warrants its own study. Programming as a discipline is broad but highly nuanced—the sociotechnical circumstance of the programming task drastically influences the nature of the programming. For example, the programming process of end-users who write software for personal use (i.e., end-user programming (Ko et al., 2011)) is fundamentally different from the programming process of data scientists who write software to explore different possibilities in code (i.e., exploratory programming (Kery and Myers, 2017; Sheil, 1986)). The programming task varies the developer’s tools (e.g., visual programming languages vs. Jupyter notebooks), focus in process (e.g., accomplishing a goal vs. exploring ideas), and key challenges (e.g., finding the right abstractions vs. managing code versions) (Ko et al., 2011; Kery and Myers, 2017). This creates disparate development experiences and necessitates the development of support tools unique to each context. Therefore, better understanding the nature of prompt programs could better inform the development of tools and environments to support the developers engaged in this type of programming activity.
We develop an understanding of prompt programming with FMs by employing the Straussian grounded theory methodology (Corbin and Strauss, 2015). This is a qualitative research methodology aimed at developing a novel exploratory explanation of a domain, referred to in such literature as a “theory”. We interview a diverse set of 20 developers who are engaged in developing prompt-powered software.
@@
This definition includes developing prompts for a range of applications like GPTs and chatbots. It also includes language agents, which are an emerging paradigm where LLMs have access to tools that they can use to complete complex tasks. This includes tools for mathematical reasoning (Schick et al., 2024) and code interpreters to execute generated code (Gao et al., 2023). However, we note that our definition of prompt programming excludes cases where a developer converses with an FM once to achieve a task (e.g., working with ChatGPT to debug the current error message in the console), as these prompts do not accept variable inputs and are not executed within a software application or code.
Through our investigation, we contribute 14 main observations (see Figure 1). While some of our results corroborate the results from Dolata et al. (2024)’s and Parnin et al. (2023)’s studies (e.g., fault localization not being certain (#11)), our results differ in notable ways. First, we find prompt programmers develop mental models (Staggers and Norcio, 1993) of the FM’s behavior on the prompt (#1) and its unique qualities (#4) by interacting with the model. While literature suggests that experts have well-formed mental models compared to novices (Staggers and Norcio, 1993), prompt programmers struggle to develop reliable mental models (#2) even after writing dozens of prompts, each with many iterations. This creates a rapid and unsystematic development process (#10). Additionally, we find that prompts can be composed and decomposed (#9) and that prompt programs’ testing occurs at different scopes (#14). While these observations exhibit some similarities to software development, the stochastic nature of FMs significantly complicates this process (see Section 5). Taken together, our 14 observations indicate that prompt programming is significantly different from traditional software development, necessitating the creation of new tools and processes to support prompt programming. These findings have implications on software practitioners, academics, tool creators, and educators interested in the role of generative AI in software engineering.
@@ 2. Related Work
We discuss related work on prompt engineering (see Section 2.1 and human aspects of software engineering for AI (see Section 2.2). Due to the quickly evolving landscape of prompt development and FMs, our work is up-to-date as of September 2024.
@@ 2.1. Prompt Engineering
Prior work has conducted empirical studies of prompt engineering. Some works have focused on chatbot development with end users through prompts (Zamfirescu-Pereira et al., 2023a, b, c). In a user study of 10 people who used a tool to create chatbots through prompts, Zamfirescu-Pereira et al. (2023c) identified several challenges of prompt development, including opportunistic prompt design approaches, a lack of systematic testing, and writing prompts that did not generalize. Jiang et al. (2022) evaluated a prompt development tool on 11 users from diverse roles, including designers, content strategist, and front-end developers. Problems participants faced included the prompt easily breaking and fixating on examples as well as having difficulty evaluating large amounts of text.
@@
Other work has investigated creating tools to facilitate prompting across a variety of contexts. This includes the creation of a prompt pattern catalogue to solve common problems encountered when conversing with a language model (LLM) (White et al., 2023), such as question refinement. A separate catalogue was created for software engineering tasks (White et al., 2024), such as specification disambiguation. Other tools include interfaces for prompting, such as PromptAid (Mishra et al., 2023), PromptMaker (Jiang et al., 2022), and PromptIDE (Strobelt et al., 2022). These tools included separate views to view the dataset, iterate the prompt, track prompt performance, and search for prompts.
@@
Perhaps most related to our study are studies that investigate prompt engineering in the context of software engineering. Dolata et al. (2024) conducted a study with 52 freelance developers on their experience with developing prompt-powered software. The participants enumerated a number of challenges, including having difficulty identifying the source of incorrect responses, budget constraints, and unrealistic client expectations. Meanwhile, Parnin et al. (2023) ran an interview study with 12 professional programmers who developed product copilots. They found that participants followed a general process of exploration, implementation, evaluation, and productization. Noted challenges included the trial and error nature of prompt development and creating benchmarks.
@@
This body of literature provides an initial understanding of the unique aspects of prompt programming. However, they are constrained to specific types of developers (i.e., end-users, professional programmers, and freelancers), limiting the generalizability of the results. Additionally, these studies do not specifically study the process of writing prompt programs with respect to software development activities. Our study builds upon these works by employing a more systematic and rigorous qualitative methodology, Straussian grounded theory (Corbin and Strauss, 2015), to study the process of writing prompt programs by intentionally sampling a broad range of prompt programmers along various axes, such as prompt complexity, programming context, and role.
2.2. Human Aspects of Software Engineering for AI
@@
Numerous works have studied the human aspects of building ML-enabled systems. This domain offers some preliminary insights on the challenges of developing prompt programs with FMs. This form of software development also involves working with and wrangling non-deterministic, opaque neural models (Giray, 2021). This is similar to prompt programming, as developers must also contend with nontransparent, stochastic FMs. Prior work describes the development of models as highly experimental (Giray, 2021; Nahar et al., 2022, 2023; Wan et al., 2019; Wolf and Paine, 2020). Literature also stresses the importance of collecting high-quality data for the domain (Nahar et al., 2022; Giray, 2021; Dolata et al., 2024; Arpteg et al., 2018; Qian et al., 2024; Nahar et al., 2023) as well as the use of quantitative metrics, like accuracy, precision, and recall, to measure model performance (Wan et al., 2019; Nahar et al., 2022).
@@
However, developing ML components introduces unique challenges for developers. Models can be difficult to debug due to its non-determinism (Giray, 2021). Multiple sources have identified the issue of training-serving skew, when training data does not generalize to production data (Nahar et al., 2022, 2023; Wan et al., 2019). To reduce this gap, models are re-trained with new data over time (Nahar et al., 2023). Additionally, expertise across a variety domains, such as software engineering and data science, are distributed across roles (Nahar et al., 2022; Wan et al., 2019; Giray, 2021). In an interview study with 45 practitioners, Nahar et al. (2022) found this could introduce collaboration challenges between different roles, such as unclear model requirements, handling evolving data, and inadequate datasets. Finally, several works note the lack of AI literacy can make requirements elicitation, communication, and collaboration with clients challenging (Nahar et al., 2022; Dolata and Crowston, 2023; Wolf and Paine, 2020).
@@
This body of literature serves as a foundation for understanding prompt programming. We extend this body of work by comparing building ML-enabled systems to prompt programming and understanding what aspects of this prior literature applies to prompt programming.
@@ 3. Methodology
Since there is limited literature on how developers create prompt programs, we use a qualitative approach to develop our understanding by exploring the process of developing prompt-powered software. In particular, we use the grounded theory methodology, which allows qualitative researchers to develop a novel exploratory explanation (called a “theory”) for a particular domain. There are three popular approaches in grounded theory: Glaserian/classical (Glaser and Strauss, 2017), Straussian (Corbin and Strauss, 2015), and constructivist (Charmaz, 2006) methodologies, which vary in their procedures and epistemology (Cole and Gillies, 2022). We present our results in Section 4 and Section 5.
Our approach is based on Straussian rounded theory (Corbin and Strauss, 2015), which involves defining a research question, developing theoretical sensitivity towards the phenomenon via a broad literature review, generating a theory through simultaneous data collection and analysis, and performing a focused literature review to contextualize the results (Cole and Gillies, 2022; Corbin and Strauss, 2015). In this section, we first discuss our grounded theory process (Section 3.1). We then describe our participants (Setion 3.2) and interview protocol (Section 3.3), and close with a discussion of the limitations to our method (Section 3.4).
@@ 3.1. Grounded Theory Process
Below, we describe our grounded theory process for the study. It contains five main stages: defining a research question, developing theoretical sensitivity, generating the grounded theory, triangulating with literature, and validating the theory. We describe this process in further detail below.
@@
Our grounded theory process began with defining a research question. We study the phenomenon where developers write a program using natural language, rather than pure code: How do programmers develop programs that are natural language prompts? Prompts can be used in a variety of contexts, such as interacting with ChatGPT in conversation (Chavan et al., 2024), but are not programs. Thus, we apply the definition of a prompt program described in Section 1.
@@
Before the study, the researchers should develop an intuition of what is occurring in the data towards the phenomenon being studied, known as theoretical sensitivity (Cole and Gillies, 2022). This can be done through professional experience and literature reviews so the researcher can extract insights from the data; however, the researcher should not be constrained by prior knowledge while developing the theory (Corbin and Strauss, 2015).
We entered the study with some theoretical sensitivity since two authors are software engineering and AI researchers and have developed multiple prompt programs themselves. Thus, following prior work (McGhee et al., 2007), we consider our background and adapt our literature review strategy to be a broad but lightweight review of prompting. The first author reviewed foundational prompting papers in machine learning venues (Brown, 2020; Wei et al., 2022) and a notable survey paper on prompting techniques in natural language processing (NLP) (Liu et al., 2023); in addition, that researcher reviewed two empirical studies on how people develop prompts (Zamfirescu-Pereira et al., 2023c; Dolata et al., 2024) from human-computer interaction (HCI) and software engineering venues. We used this literature review in combination with the authors’ existing domain expertise to develop questions for the interview protocol (see Section 3.3).
@@
We conducted 60-minute interviews with a diverse set of 20 developers who had developed prompt programs (see Table 1) over Zoom. Interviews were audio recorded and transcribed; recordings were deleted afterwards. Two authors were present for the first 6 interviews to become familiar with the data; the first author conducted the remaining interviews.
To develop the initial theory, three authors first independently performed line-by-line coding on the first two interviews in separate codebooks to become further sensitized to the data. Each code contained a description of the code as well as observations from the interviews. They then reconvened to merge the individual codebooks by identifying codes with similar concepts and merging them into a shared codebook. The remaining codes were then discussed and added or removed to the codebook based on unanimous vote. The first author then coded the next interview; two authors reviewed the codes for agreement. We identified five instances of disagreement. These disagreements were then discussed and resolved. Next, the three authors performed axial coding, grouping the emerging codes into preliminary categories upon unanimous vote.
For the remaining interviews, the authors individually open coded the interview transcripts, making note of any new codes that emerged in the shared codebook. Following the open coding, a memo for each interview was created to record notable parts of the interview to further refine the emerging theory. The authors met regularly to discuss the new codes and observations. New codes and observations were added to the shared codebook upon unanimous vote and the categories were further refined as more data was collected.
As the theory developed, we performed maximum variation sampling (Suri, 2011) to obtain a diverse set of participants that could challenge or extend the theory. We recruited participants through snowball sampling within the authors’ social networks and in online open-source communities who met the definition of creating a prompt program (see Section 1). Participants were recruited based on the types of model used (e.g., open-source models vs. closed-source models, vision language models vs. language models), application domain (e.g., robotics, education, productivity), organization size (e.g., large technology company vs. startup), programming context (e.g., academia, industry, freelance, open-source software), role, and/or prompt complexities (e.g., a language agent that is able to use tools vs. single prompt vs. multi-prompt). We initially achieved theoretical saturation after interviewing 14 participants. We continued to recruit participants to ensure our observations held across a variety of situations, including in open-source software and startups. We stopped data collection after reaching 20 participants.
| ID | Context | Exp. | # Prompts | Complexity | Model | App Domain | Task |
@@
While developing the theory on prompt programming, we triangulated our findings with prior literature. Since the relevant literature was not identifiable via keyword search, we follow prior work (Nahar et al., 2022) in performing a focused literature review by forward and/or backward snowballing the set of 5 papers from the initial literature review. Relevant literature spanned several communities, including NLP, HCI, programming languages, software engineering, and machine learning. Of the 502 papers considered, we identified 33 as relevant. See the supplemental materials for this set of papers (Authors, 2024). The first author applied codes from the emerging theory to the papers, making updates to the codes as necessary to further refine the theory.
@@
To validate the theory, we assess the fit of the theory. Straussian rounded theory suggests validating the theory with professionals and study participants to see if they resonate with the findings (Corbin and Strauss, 2015). To this end, we e-mailed the participants and an external researcher who studies prompt programming a summary of the theory as well as a draft of Section 4 and Table 1. We asked them to evaluate the correctness of the theory as well as any feedback for the theory based on their personal experiences. In total, 8 people responded to this inquiry. All respondents indicated general agreement with the findings, some with a few wording clarifications.
Additionally, to assess the coding reliability between authors, the first author randomly selected and open coded two interviews at the category level. The labels were then removed with the original highlighted text spans remaining intact. One author was assigned to each interview and applied the code categories to each highlighted text span. Based on this procedure, our inter-rater reliability was 0.89 and 0.81 based on Cohen’s —both almost perfect agreements (Landis and Koch, 1977).

### v2

@@ Abstract.
Generative pre-trained models power intelligent software features used by millions of users controlled by developer-written natural language prompts. Despite the impact of prompt-powered software, little is known about its development process and its relationship to programming. In this work, we argue that some prompts are programs and that the development of prompts is a distinct phenomenon in programming known as “prompt programming”. We develop an understanding of prompt programming using Straussian grounded theory through interviews with 20 developers engaged in prompt development across a variety of contexts, models, domains, and prompt structures. We contribute 15 observations to form a preliminary understanding of current prompt programming practices. For example, rather than building mental models of code, prompt programmers develop mental models of the foundation model (FM)’s behavior on the prompt by interacting with the FM. While prior research shows that experts have well-formed mental models, we find that prompt programmers who have developed dozens of prompts still struggle to develop reliable mental models. Our observations show that prompt programming differs from traditional software development, motivating the creation of prompt programming tools and providing implications for software engineering stakeholders.
@@ 1. Introduction
Generative pre-trained models (e.g., GPT-4 (Achiam et al., 2023), Dall-E (Betker et al., 2023))—also known as foundation models (FMs)—have changed how programmers build software. AI programming assistants that generate code (e.g., GitHub Copilot (Chen et al., 2021)) have improved developer productivity (Ziegler et al., 2022; Peng et al., 2023; Ziegler et al., 2024) by helping developers write significant portions of code, learn new APIs and programming languages, and write tests (Liang et al., 2024). Recently, instruction-tuned FMs (Ouyang et al., 2022) like ChatGPT (ope, 2024a) have expanded the assistance with software development tasks by writing natural language prompts. This includes resolving code issues, developing new features, refactoring, and information seeking (Chavan et al., 2024; Haque et al., 2024).
@@
The rising prominence of prompts has ushered in “prompt engineering”, whereby FM users repeatedly write and revise natural language prompts. These prompts enable new intelligent features when integrated in popular software applications (Parnin et al., 2023), such as Google Search (goo, 2024) and Microsoft Office (mic, 2024b), reaching millions of users (mic, 2024a). As of January 2024, engineered prompts have also powered over 3 million custom versions of ChatGPT for specific tasks, known as GPTs (ope, 2024b).
@@
Yet, little is known about prompt engineering and its relationship to programming. Relevant works include Dolata et al. (2024)’s study of 52 freelance developers building solutions based on generative AI and Parnin et al. (2023)’s study of 26 professional software developers integrating generative AI into products. While these studies offer insight on prompt development, they are constrained to freelance and professional software development, limiting the generalizability of the results to other programming contexts. We address these gaps by following a more systematic and rigorous qualitative methodology—Straussian grounded theory (Corbin and Strauss, 2015)—to study a diverse sample of programmers to understand the process of developing prompts embedded in software applications.
In this work, we argue that some prompts function as programs, a phenomenon we call prompt programming (Jiang et al., 2022). In prompt programming, prompts are authored at design-time and executed at runtime with variable user inputs, but are written in natural language instead of a programming language. Prompt programming is a specific form of prompt engineering—broadly defined as iterating on prompts to improve FM outputs (Zamfirescu-Pereira et al., 2023c)—where the same prompt handles variable inputs given at runtime. For example, a prompt program on a teleconferencing platform may take a meeting transcript and attendee name and return bulleted todos and deadlines for that individual.
In prompt programming, natural language is a programming abstraction, but unlike formalisms such as DSLs or machine code, it introduces ambiguity and reduces precision (Dijkstra, 1979). While the idea of programming in natural language dates as early as the 1970s, where Dijkstra discussed a vision of instructing machines in “our native tongues” (Dijkstra, 1979), it has become feasible and widespread with FMs due to instruction-tuning (Ouyang et al., 2022) that teaches models to follow natural language instructions.
@@
We argue that prompt programming is a phenomenon that warrants its own study, as the sociotechnical circumstance of the programming task influences the nature of the programming. For example, the programming process of end-users who write software for personal use (i.e., end-user programming (Ko et al., 2011)) is different from that of data scientists who write software to explore different possibilities in code (i.e., exploratory programming (Kery and Myers, 2017; Sheil, 1986)), and both differ from a professional programmer who focuses on a small component within a larger system. The programming task also varies with the developer’s tools (e.g., visual programming languages vs. Jupyter notebooks), focus on the formality of the process (e.g., accomplishing a goal vs. exploring ideas), and key challenges (e.g., finding the right abstractions vs. managing code versions) (Ko et al., 2011; Kery and Myers, 2017). This creates disparate experiences and necessitates the development of support tools unique to each context. Thus, studying prompt programs could inform the development of tools to support prompt programmers.
@@
This definition includes developing prompts for applications like GPTs and chatbots. It also covers language agents, where large language models (LLM) receive arbitrary user requests and execute them in dynamic environments using external modules known as “tools” (e.g., modules for mathematical reasoning (Schick et al., 2024) and code interpreters (Gao et al., 2023)). However, this definition excludes cases where a developer converses with an FM once to achieve a task (e.g., prompt engineering with ChatGPT to debug a specific error message). These prompts do not accept variable inputs, do not generalize beyond the specific case, and are not executed within a broader software application.
[…]
@@ 2. Related Work
We discuss related work on prompt engineering (Section 2.1) and software engineering for AI (SE4AI; see Section 2.2). Due to the rapidly evolving landscape of prompt engineering and FMs, this discussion will necessarily miss more recent developments.
@@ 2.1. Prompt Engineering
Prior literature has examined prompt engineering in a variety of contexts. This includes user studies on prompt engineering, prompt engineering tools, and retrospective interview studies on prompt engineering in software development. We discuss this in more detail below.
@@
Previous work has conducted empirical user studies of prompt engineering. Some works have focused on the development of chatbots with end users using prompts (Zamfirescu-Pereira et al., 2023a, b, c). In a user study of 10 people who used a tool to create chatbots through prompts, Zamfirescu-Pereira et al. (2023c) identified several challenges of prompt development, including opportunistic prompt design approaches, a lack of systematic testing, and writing prompts that were not generalized. Jiang et al. (2022) evaluated a prompt development tool in 11 users in various roles, including designers, content strategists, and front-end developers. The problems participants faced included the prompt easily breaking and overfitting on examples, as well as having difficulty evaluating large amounts of text.
@@
Other work has investigated the creation of prompting tools to assist with various prompting tasks. Prompt pattern catalogs can solve common problems encountered when conversing with an LLM (White et al., 2023), such as question refinement, as well as software engineering tasks (White et al., 2024), such as specification disambiguation. Other relevant tools include prompting interfaces, such as PromptAid (Mishra et al., 2025), PromptMaker (Jiang et al., 2022), and PromptIDE (Strobelt et al., 2022). These tools included separate views to show the dataset, iterate on the prompt, track the performance of the prompt, and search for the prompts.
@@
Most related to this study are retrospective interview studies that investigate prompt development in software engineering. Dolata et al. (2024) conducted a study with 52 freelance developers on their experience with building prompt-powered software. Their interview focuses on the positive and negative experiences with generative AI, project uncertainties, and views on freelancing. The methodology involved recruiting from Upwork, conducting retrospective interviews, performing a thematic analysis, and comparing the findings with four meta-reviews of SE4AI literature. The participants enumerated several challenges, including having difficulty identifying the source of incorrect responses, budget constraints, and unrealistic client expectations. Meanwhile, Parnin et al. (2023) conducted an interview study with 26 professional programmers who developed product copilots. Interview topics included motivation for using AI in products, major tasks for building the generative AI application, prompt engineering, testing, tooling, challenges, learning-related skills, and concerns with AI. Their methodology involved recruiting participants from UserInterviews.com, conducting retrospective interviews and structured brainstorming sessions, and performing a thematic analysis. They found that the participants followed a general process of exploration, implementation, evaluation, and productization. This work noted several challenges, including the trial-and-error nature of prompt development and creating benchmarks.
@@
This literature provides an initial understanding of the unique aspects of prompt programming. However, they are constrained to specific types of developers (i.e., end-users, professional programmers, and freelancers), limiting the generalizability of the results. Additionally, these works do not study the process of writing prompt programs with respect to software development activities. Our study builds upon these works by employing a more systematic and rigorous qualitative methodology, Straussian grounded theory (Corbin and Strauss, 2015), to study the process of writing prompt programs. We sample a broad range of prompt programmers along various axes, such as prompt structure, programming context, and role, as well as ask questions about the end-to-end development process.
2.2. Software Engineering for AI
Numerous works have studied the human aspects of building ML-enabled systems, also known as software engineering for AI. These works have documented the experiences of practitioners and their common challenges in this domain. We elaborate further below.
@@ 2.2. Software Engineering for AI
Building ML-enabled systems involves working with and wrangling nondeterministic, opaque neural models (Giray, 2021). This is similar to prompt programming, as developers must also contend with non-transparent, stochastic FMs. Previous work describes the development of models as highly experimental (Giray, 2021; Nahar et al., 2022, 2023; Wan et al., 2019; Wolf and Paine, 2020). The literature also stresses the importance of collecting high-quality data for the domain (Nahar et al., 2022; Giray, 2021; Dolata et al., 2024; Arpteg et al., 2018; Qian et al., 2024; Nahar et al., 2023) as well as the use of quantitative metrics, such as precision, precision, and recall, to measure model performance (Wan et al., 2019; Nahar et al., 2022).
@@
However, the development of ML components presents unique challenges. Models can be difficult to debug due to their determinism (Giray, 2021), while training-serving skew can arise when training data fail to generalize to production (Nahar et al., 2022, 2023; Wan et al., 2019). To mitigate this, models are periodically re-trained (Nahar et al., 2023). In addition, expertise in a variety of domains, such as software engineering and data science, is distributed between roles (Nahar et al., 2022; Wan et al., 2019; Giray, 2021). In an interview study with 45 practitioners, Nahar et al. (2022) found this introduced collaboration challenges between different roles, such as unclear model requirements, handling evolving data, and inadequate datasets. Finally, a lack of AI literacy can make requirements elicitation, communication, and collaboration with clients challenging (Nahar et al., 2022; Dolata and Crowston, 2023; Wolf and Paine, 2020).
@@
This body of literature serves as a foundation for understanding prompt programming. We extend this body of work by comparing the building of ML-enabled systems to prompt programming and understanding what aspects of this prior literature apply to prompt programming.
@@ 3. Methodology
Since there is limited literature on prompt programming, we used a qualitative approach to explore the process of developing prompt-powered software. We used the grounded theory methodology, which allows researchers to develop a novel exploratory explanation for a domain.
There are three popular approaches to grounded theory: Glaserian/classical (Glaser and Strauss, 2017), Straussian (Corbin and Strauss, 2015), and constructivist (Charmaz, 2006) methodologies, which vary in their procedures and epistemology (Cole and Gillies, 2022). Our approach is based on Straussian grounded theory (Corbin and Strauss, 2015) (see Figure 2). This involves defining a research question; developing theoretical sensitivity to the phenomenon through prior literature and domain expertise; generating a grounded theory via diverse sampling and simultaneous data collection and analysis; and validating the grounded theory by triangulating with literature, performing an assessment of the fit of the grounded theory, and computing inter-rater reliability (Cole and Gillies, 2022; Corbin and Strauss, 2015). In this section, we first discuss our grounded theory process (Section 3.1). We then describe our participants (Section 3.2) and interview protocol (Section 3.3), and close with a discussion of the limitations of our method (Section 3.4). We present our results in Section 4 and Section 5.
@@ 3.1. Grounded Theory Process
We describe our grounded theory process below. It contains five main stages: defining a research question, developing theoretical sensitivity, generating the grounded theory, triangulating with the literature, and validating the grounded theory. We describe this process in further detail below.
@@
Our process began with defining a research question. We study the phenomenon where developers write a prompt program using natural language prompts, rather than pure code: How do programmers develop programs that incorporate natural language prompts? Prompts can be used in a variety of contexts, such as interacting with ChatGPT in conversation (Chavan et al., 2024), but are not programs. Thus, we apply the definition of a prompt program described in Section 1.
@@
Before the study, the investigator should develop an intuition for the phenomenon being studied, known as theoretical sensitivity (Cole and Gillies, 2022). This can be done through professional experience and understanding the prior literature so that researchers can extract insights from the data. However, the investigators should not be constrained by prior knowledge while developing the grounded theory (Corbin and Strauss, 2015). We entered the study with some theoretical sensitivity since two authors are software engineering and AI researchers and have developed multiple prompt programs. Following prior work (McGhee et al., 2007), we considered our background and opted for a lightweight overview of prompting literature. We considered foundational prompting papers in machine learning venues (Brown et al., 2020; Wei et al., 2022), a notable survey paper on prompting techniques in natural language processing (NLP) (Liu et al., 2023) as well as two empirical studies on how people develop prompts (Zamfirescu-Pereira et al., 2023c; Dolata et al., 2024) from human-computer interaction (HCI) and software engineering venues. We considered this literature and the authors’ existing domain expertise to develop the interview protocol (Section 3.3).
@@
We conducted 45-minute interviews with a diverse set of 20 prompt programmers (see Table 1) over Zoom. The interviews were recorded and transcribed; the recordings were later deleted. Two authors were present for the first six interviews to become familiar with the data. The first author conducted the remaining interviews.
To develop the initial grounded theory, three authors first independently performed line-by-line coding on the first two interviews in separate codebooks to become more sensitized to the data. Each code contained a description of the code as well as observations from the interviews. The authors then reconvened to merge the individual codebooks by identifying codes with similar concepts and merging them into a shared codebook. The remaining codes were then discussed and added or removed to the codebook by unanimous vote. The first author then coded the next interview, and the other authors reviewed the codes for agreement. We identified five instances of disagreement, which were discussed and resolved. The three authors then performed axial coding, grouping the emerging codes into preliminary categories on unanimous vote.
For the remaining interviews, the authors individually open coded the interview transcripts, noting new codes that emerged in the shared codebook. A memo was created for each interview to capture key insights and refine the emerging grounded theory. The authors met regularly to discuss the new codes and observations. New codes and observations were added to the shared codebook upon unanimous vote, and the categories were further refined as more data was collected.
As the grounded theory developed, we performed maximum variation sampling (Suri, 2011) to obtain a diverse set of participants that could challenge or extend the grounded theory. We recruited participants who met the definition of creating a prompt program (Section 1) by snowball sampling within the authors’ social networks and recruiting in online open-source communities. Participants were recruited based on the types of models used (e.g., open-source models vs. closed-source models, vision language models vs. language models), application domain (e.g., robotics, education, productivity), organization size (e.g., large technology company vs. startup), programming context (e.g., academia, industry, freelance, open-source software), role, and/or prompt structure (e.g., a language agent that is able to use tools vs. single prompt vs. multi-prompt). We initially achieved theoretical saturation after interviewing 14 participants. We continued to recruit participants to ensure that our observations held across a variety of situations, including open-source software and startups. We stopped data collection after reaching 20 participants.
| ID | Context | Exp. | # Prompts | Structure | Model | App Domain | Task |
@@
While developing the grounded theory of prompt programming, we triangulated our findings with the prior literature. Since the relevant literature was not identifiable through keyword search, we followed previous work (Nahar et al., 2022) in performing a focused, best-effort literature review by performing forward and/or backward snowballing on the initial set of five papers. The relevant papers were first identified based on titles related to building AI systems or prompt engineering. They were later verified by reading the manuscript. In the end, this literature spanned several communities, including NLP, HCI, programming languages, software engineering, and machine learning. Of the 502 papers considered, we identified 33 as relevant. See the supplemental materials for this set of papers (Liang et al., 2025). For each paper, the first author identified text related to prompt programming and applied codes from the codebook, updating the codes as necessary to further refine the grounded theory.
@@
To validate the grounded theory, we assess its fit. Straussian grounded theory suggests validating the grounded theory with professionals and study participants to see if they resonate with the findings (Corbin and Strauss, 2015). To this end, we sent an email containing a summary of the grounded theory as well as a draft of Section 4, Section 5, and Table 1 to the 20 participants and one external researcher who studies prompt programming. We asked them to reflect on the fit of the grounded theory and provide feedback based on their experience. Nine people responded to this inquiry and indicated general agreement with the findings, some with wording changes.
Additionally, to assess the reliability of the coding, the first author randomly selected and open-coded two interviews at the category level. The labels were then removed, with the original highlighted text spans remaining intact. One author was assigned to each interview and applied the code categories to each highlighted text span. Based on this procedure, our inter-rater reliability was 0.89 and 0.81 based on Cohen’s —indicating that both are almost in perfect agreement (Landis and Koch, 1977).